10. Potassium hydrogen phthalate, abbreviated KHP, is used to standardize NaOH A KHP solutions. standard KHP solution is made by dissolving 2.12 grams of KHP in 100.00 mL of the endpoint. solution is then titrated with NaoH solution. It takes 23.12 mL of NaoH to reach What is the concentration of the NaoH?
Answer #1

molar mass of KHP = 204.22 gm/mol then 2.12 gm of KHP = 2.12 / 204.22 = 0.010381 mole

Neutrilization reaction between KHP and NaOH take place in equimolar proportion thus to react with 0.010381 mole of KHP required NaOH = 0.010381 mole

23.12 ml = 0.02312 L

Molarity = no. of mole / volume of solution in liter

Molarity of NaOH = 0.010381 / 0.02312 = 0.449 M

Concentration of NaOH soution = 0.449 M